Can I please get some help with this question?
Interest Rate. A credit manager rates each applicant for a car loan on a scale of 1 through 5 and then determines the interst rate from the accompanying table. Find the equation of the line in slope intercept form that goes through these points.
Credit Rating 1 2 3 4 5
Interest 24 20 16 12 8
Rate%

Answer by venugopalramana(3286)

Credit Rating=R.... 1...... 2...... 3....... 4...... 5
Interest ....=I.....24..... 20..... 16...... 12..... 8
CHANGE IN I= 20-24=-4
CORRESPONDING CHANGE IN R=2-1=1
RATIO=-4/1=-4
TEST OTHER POINTS.....(16-20)/(3-2)=(12-16)/(4-3)=(8-12)/(5-4)=-4
SO THE EQN.IS
I-24=-4(R-1)
I=-4R+28